### v2

# Upper bound for the total mean curvature of spin fill-ins

Abstract.
Gromov conjectured that the total mean curvature of the boundary of a compact Riemannian manifold can be estimated from above by a constant depending only on the boundary metric and on a lower bound for the scalar curvature of the fill-in. We prove Gromov’s conjecture if the manifolds are spin and the mean curvature $H$ is non-negative. One can also allow $H$ to take negative values, but then the constant depends on a (negative) lower bound for $H$.
Key words and phrases:
Spin fill-in, mean curvature bound, scalar curvature, Dirac operator
2020 Mathematics Subject Classification:
53C20, 53C27
1. Introduction
Given a compact Riemannian spin manifold $M$ without boundary and a number $\lambda\geq 0$, we may consider all compact connected manifolds $X$ that bound $M$ as a Riemannian spin manifold and satisfy the scalar curvature bound $\mathrm{scal}_{X}\geq-\lambda^{2}$. Given such an $X$, let $H$ be the unnormalized mean curvature of the boundary $\partial X=M$. Our sign convention for the mean curvature is such that the boundary of an $(n+1)$-dimensional Euclidean ball of radius $r$ has mean curvature $H=n/r$.
Of course, $M$ may not bound any compact manifold $X$, but if it does so, it follows from [SWW]*Theorem 1.1 by Shi, Wang, and Wei that the given metric on $M$ can be extended to a metric on $X$ with $\mathrm{scal}_{X}>0$. Their result gives no control on the mean curvature of the boundary. Theorem 3.7 in [BH] by the author and Hanke implies that the metric on $X$ can be deformed in such a way that its scalar curvature remains positive, it still induces the given metric on $M$, and the mean curvature of the boundary becomes arbitrarily negative. Thus
$\inf\int_{M}H=$-$\infty$
where the infimum is taken over $X$ with $\mathrm{scal}_{X}>0$ and $\partial X=M$ (as Riemannian manifolds).
Gromov conjectured that $\int_{M}H$ cannot be made arbitrarily large. More precisely, the following is believed to hold:
Conjecture (Gromov [G]*p. 232).
Let $M$ be a compact Riemannian manifold without boundary which bounds a compact manifold and let $\sigma\in\mathbb{R}$. Then there exists a constant $C(M,\sigma)$ such that for each compact Riemannian manifold $X$ with boundary $\partial X=M$ and scalar curvature $\mathrm{scal}_{X}\geq\sigma$ we have
$\int_{M}H\leq C(M,\sigma).$
If $\dim(M)=1$ and $\sigma=0$, then this follows from the Gauss-Bonnet theorem. Indeed, assuming without loss of generality that $X$ is connected, the Gauss-Bonnet theorem gives us
$\int_{M}H=2\pi\chi(X)-\frac{1}{2}\int_{X}\mathrm{scal}_{X}\leq 2\pi.$
If we are more modest and replace the integral of the mean curvature by its minimum $H_{\min}=\min_{M}H$, then we indeed have
$H_{\min}\leq C(M,\sigma)$ (1)
for spin manifolds. This follows from [HMR]*Theorem 2 by Hijazi, Montiel, and Roldán. Here the constant $C(M,\sigma)$ depends on the first non-negative eigenvalue of the Dirac operator of $M$ and on $\sigma$. Combining this with an upper Dirac eigenvalue estimate by the author ([Baer]*Main Theorem), one obtains (1) where the constant depends on the hyperspherical radius of $M$ and on $\sigma$, as has been observed by Brendle, Tsiamis, and Wang in [BTW]. See also [CHZ]*Theorem 1.5 by Cecchini, Hirsch, and Zeidler for (1) in the case $\sigma=0$. Clearly, estimate (1) is only meaningful if the mean curvature is positive.
We prove Gromov’s conjecture for spin manifolds if the mean curvature is non-negative. More precisely, we show the following:
Theorem.
Let $X$ be a compact Riemannian spin manifold of dimension $n+1\geq 2$ with smooth boundary $\partial X=M$. Let $\lambda\geq 0$ be such that the scalar curvature of $X$ satisfies $\mathrm{scal}_{X}\geq-\lambda^{2}$. Let $\eta\geq 0$ be such that the mean curvature $H$ of the boundary satisfies $H\geq-\eta$. Then we have
$\int_{M}H\leq C(M)+\Big(\sqrt{\tfrac{n}{n+1}}\,\lambda+\eta\Big)\mathsf{vol}(M),$
where $C(M)$ is a constant depending only on the Riemannian spin manifold $M$.
All previously known results need the assumption that $H\geq 0$ or even $H>0$. If $\lambda=0$, $H>0$, $X$ is spin, and each connected component of $M$ admits an isometric embedding as a strictly convex hypersurface in $\mathbb{R}^{n+1}$, then this has been proved by Shi and Tam in [ST]*Theorems 1 and 4.1. The constant $C(M)$ is in this case given by the total mean curvature of the isometric embedding of $M$ into $\mathbb{R}^{n+1}$. The condition that each connected component of $M$ embeds isometrically as a strictly convex hypersurface into $\mathbb{R}^{n+1}$ has been relaxed by Eichmair, Miao, and Wang in [EMW] to require the components to have non-negative scalar curvature and admit embeddings as star-shaped hypersurfaces.
The case where $M$ is diffeomorphic to a sphere with $\lambda=0$, $H>0$, and $X$ spin has been treated by Shi, Wang, and Wei in [SWW]*Theorem 4.1.
Mantoulidis and Miao investigated in [MM] the case where $n=2$ and $M$ has arbitraty topology, again if $\lambda=0$ and $H>0$.
If $M=T^{n}$ is a torus, $X=D^{2}\times T^{n-1}$, $n\leq 6$, and $H>0$, then Gromov’s conjecture has been proved by Wang in [W]*Theorems 1.6 and 1.8. In this case, $\lambda$ is allowed to be negative. If the torus $M$ carries a flat metric, then $C(M)$ is related to the systole of $M$.
In the mean time, Frenck, Hanke, and Hirsch have presented a different proof of Gromov’s conjecture if $H\geq 0$, using surgery and deformation methods ([FHH]). Their approach has the advantage that it also works in the non-spin case if $n\leq 6$.
Acknowledgments: This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – project ID 569831821.
2. A boundary value problem
Let $X$ be a compact connected Riemannian spin manifold of dimension $n+1\geq 2$ with boundary $\partial X=M$. The spinor bundle $\Sigma X$ of $X$ is a Hermitian vector bundle with metric connection. We choose the convention that the fiberwise scalar product of $\Sigma X$ is antilinear in the first argument and linear in the second argument. Let $\gamma\colon TX\to\operatorname{\mathsf{End}}(\Sigma X)$ denote the Clifford multiplication. It satisfies the Clifford relations
$\gamma(v)\gamma(w)+\gamma(w)\gamma(v)=-\langle v,w\rangle\operatorname{\mathrm{id}}_{\Sigma X}$
for all tangent vectors $v,w\in T_{p}X$ and $p\in X$. The Dirac operator of $X$ is denoted by $D$.
Denote the inward unit normal vector field along the boundary by $\nu$. Then
$s:=i\gamma(\nu)\in\operatorname{\mathsf{End}}(\Sigma X)$
is a self-adjoint involution. It anticommutes with Clifford multiplication by tangent vectors of $M$. Hence, the eigenspaces corresponding to the eigenvalues $\pm 1$ of $s$ have the same dimension. We obtain an orthogonal vector bundle decomposition
$\Sigma X|_{M}=\Sigma^{+}M\oplus\Sigma^{-}M,$ (2)
where $s$ acts on $\Sigma^{\pm}M$ as $\pm 1$. We denote the orthogonal projections onto these subbundles by
$P^{\pm}\colon\Sigma X|_{M}\to\Sigma^{\pm}M.$
If $n$ is even, then $\Sigma X|_{M}$ can be canonically identified with the spinor bundle $\Sigma M$ of $M$. If $n$ is odd, then $\Sigma X|_{M}$ can be canonically identified with $\Sigma M\oplus\Sigma M$ (which is not the same decomposition as the splitting in (2)). Denoting the intrinsic Dirac operator of $M$ by $\tilde{D}_{M}$ and using these identifications, we set $D_{M}:=\tilde{D}_{M}$ if $n$ is even and $D_{M}=\begin{pmatrix}\tilde{D}_{M}&0\\ 0&-\tilde{D}_{M}\end{pmatrix}$ if $n$ is odd. Then $D_{M}$ anticommutes with $s$ in both cases (see [Baer3]*Proposition 2.3) and hence interchanges the subbundles $\Sigma^{+}M$ and $\Sigma^{-}M$. Moreover, by [Baer3]*Proposition 2.2, we have the relation
$\nabla_{\nu}=-\gamma(\nu)D-D_{M}+\tfrac{1}{2}H.$ (3)
This implies, in particular, that $D_{M}$ is an adapted boundary operator for $D$ in the sense of [BB].
We denote the space of square integrable sections of a vector bundle $E$ over $X$ by $L^{2}(X,E)$ and the $L^{2}$-Sobolev space of order $\alpha$ by $H^{\alpha}(X,E)$ and similarly for $M$.
Lemma 1.
Let $X$ be a compact Riemannian spin manifold with boundary $\partial X=M$. Let $\Phi\in H^{1}(X,\Sigma X)$ satisfy $(D+i\lambda)\Phi=0$ where $\lambda\geq 0$. Then we have
$\|P^{-}(\Phi|_{M})\|_{L^{2}(M)}\geq\|P^{+}(\Phi|_{M})\|_{L^{2}(M)}.$
Similarly, if $\lambda\leq 0$, then
$\|P^{-}(\Phi|_{M})\|_{L^{2}(M)}\leq\|P^{+}(\Phi|_{M})\|_{L^{2}(M)}.$
Proof.
By Green’s formula for the Dirac operator (see e.g. [BB]*Lemma 2.6) we have
$0$ $=\int_{X}\langle(D+i\lambda)\Phi,\Phi\rangle-\int_{X}\langle\Phi,(D+i\lambda)\Phi\rangle$
$=\int_{X}\langle\Phi,(D-i\lambda)\Phi\rangle-\int_{M}\langle\gamma(\nu)\Phi,\Phi\rangle-\int_{X}\langle\Phi,(D+i\lambda)\Phi\rangle$
$=-\int_{M}\langle\gamma(\nu)\Phi,\Phi\rangle-2i\lambda\|\Phi\|_{L^{2}(X)}^{2}$
$=-i\int_{M}\langle s\,\Phi,\Phi\rangle-2i\lambda\|\Phi\|_{L^{2}(X)}^{2}$
$=-i\int_{M}\langle P^{+}\Phi-P^{-}\Phi,P^{+}\Phi+P^{-}\Phi\rangle-2i\lambda\|\Phi\|_{L^{2}(X)}^{2}$
$=-i\Big(\|P^{+}\Phi\|_{L^{2}(M)}^{2}-\|P^{-}\Phi\|_{L^{2}(M)}^{2}+2\lambda\|\Phi\|_{L^{2}(X)}^{2}\Big).$
This implies
$\|P^{-}\Phi\|_{L^{2}(M)}^{2}=\|P^{+}\Phi\|_{L^{2}(M)}^{2}+2\lambda\|\Phi\|_{L^{2}(X)}^{2},$ (4)
which proves the lemma. ∎
As a consequence we find the well-posedness of the following boundary value problem:
Proposition 1.
Let $X$ be a compact connected Riemannian spin manifold with boundary $\partial X=M$ and let $\lambda\geq 0$. For each $\varphi\in H^{\nicefrac{{1}}{{2}}}(M,\Sigma^{+}M)$ and $\Psi\in L^{2}(X,\Sigma X)$ there exists a unique $\Phi\in H^{1}(X,\Sigma X)$ such that
$(D-i\lambda)\Phi=\Psi\quad\text{and}\quad P^{+}(\Phi|_{M})=\varphi.$
Proof.
Both boundary conditions $P^{+}(\Phi|_{M})=0$ and $P^{-}(\Phi|_{M})=0$ are $\infty$-regular elliptic in the sense of [BB]. Therefore the four operators
$(D\pm i\lambda)\oplus P^{\pm}(\cdot|_{M})\colon H^{1}(X,\Sigma X)\to L^{2}(X,\Sigma X)\oplus H^{\nicefrac{{1}}{{2}}}(M,\Sigma^{\pm}M)$ (5)
are Fredholm. If $\Phi\in H^{1}(X,\Sigma X)$ lies in the kernel of $(D-i\lambda)\oplus P^{+}(\cdot|_{M})$, then $(D-i\lambda)\Phi=0$ and $P^{+}(\Phi|_{M})=0$. By Lemma 1, we then also have $P^{-}(\Phi|_{M})=0$. Hence, $\Phi|_{M}=0$. By the unique continuation property for Dirac-type operators, we conclude $\Phi=0$ (see e.g. the proof of [BBHW]*Corollary B.2) [^1] This also follows from (4) unless $\lambda=0$.. Thus, the operator
$(D-i\lambda)\oplus P^{+}(\cdot|_{M})\colon H^{1}(X,\Sigma X)\to L^{2}(X,\Sigma X)\oplus H^{\nicefrac{{1}}{{2}}}(M,\Sigma^{+}M)$ (6)
is injective. The adjoint boundary value problem is given by
$(D+i\lambda)\oplus P^{-}(\cdot|_{M})\colon H^{1}(X,\Sigma X)\to L^{2}(X,\Sigma X)\oplus H^{\nicefrac{{1}}{{2}}}(M,\Sigma^{-}M)$
which, by the same reasoning, is also injective. Therefore, the operator in (6) is surjective, and hence an isomorphism. ∎
3. Some preparation
For the proof of the theorem we need some technical preparation. Only Proposition 2 will be needed later on.
Lemma 2.
Let $V$ be a finite-dimensional Euclidean or unitary vector space. Let $e_{1},...,e_{n}$ be an orthonormal basis of $V$. Let $U\subset V$ be a $k$-dimensional subspace and let $P\colon V\to U$ be the orthogonal projection onto $U$. Then we have:
$\sum_{j=1}^{n}|Pe_{j}|^{2}=k.$
Proof.
We choose an orthonormal basis $u_{1},...,u_{k}$ of $U$. Then we have
$u_{m}=\sum_{j=1}^{m}\langle u_{m},e_{j}\rangle e_{j}\quad\text{ and hence }\quad|u_{m}|^{2}=\sum_{j=1}^{k}|\langle u_{m},e_{j}\rangle|^{2}.$
Similarly, we have
$|Pe_{j}|^{2}=\sum_{m=1}^{k}|\langle e_{j},u_{m}\rangle|^{2}.$
We compute
$\sum_{j=1}^{n}|Pe_{j}|^{2}$ $=\sum_{j=1}^{n}\sum_{m=1}^{k}|\langle u_{m},e_{j}\rangle|^{2}=\sum_{m=1}^{k}|u_{m}|^{2}=k.\qed$
Lemma 3.
Let $M$ be a manifold (possibly with boundary) and let $U_{j}\subset M$ be an open cover of $M$. Then there exist smooth non-negative functions $\chi_{j}\in C^{\infty}(M,\mathbb{R})$ such that $\mathrm{supp}(\chi_{j})\subset U_{j}$ for every $j$ and $\sum_{j}\chi_{j}^{2}\equiv 1$ on $M$.
Proof.
It is well known that there exists a smooth partition of unity $\psi_{j}\in C^{\infty}(M,\mathbb{R})$ subordinate to the open cover, i.e., $\psi_{j}\geq 0$ and $\mathrm{supp}(\psi_{j})\subset U_{j}$ for every $j$ and $\sum_{j}\psi_{j}\equiv 1$ on $M$. Then
$u:=\sum_{j}\psi_{j}^{2}$
is smooth and positive on $M$. The functions $\chi_{j}:=\psi_{j}/\sqrt{u}$ have the desired properties. ∎
Proposition 2.
Let $M$ be a compact manifold (possibly with boundary) and let $E\to M$ be a Hermitian or Riemannian vector bundle. Then there exist finitely many sections $\varphi_{1},...,\varphi_{m}\in C^{\infty}(M,E)$ such that for each subbundle $F\subset E$ of rank $k$ we have
$\sum_{j=1}^{m}|P\varphi_{j}|^{2}\equiv k$
where $P\colon E\to F$ is the orthogonal projection onto $F$.
Proof.
We choose a finite open cover $U_{j}$ of $M$ such that $E|_{U_{j}}$ is trivial for each j=1,,Nj=1,...,N. We further choose locally defined smooth sections $e_{j,1},...,e_{j,r}$ of $E|_{U_{j}}$ which form an orthonormal basis at each point of $U_{j}$. By Lemma 3 there exist smooth functions $\chi_{j}\in C^{\infty}(M,\mathbb{R})$ such that $\mathrm{supp}(\chi_{j})\subset U_{j}$ for every $j$ and $\sum_{j}\chi_{j}^{2}\equiv 1$ on $M$. The sections $\chi_{j}\cdot e_{j,l}$ can be extended by zero to smooth sections in $C^{\infty}(M,E)$. We define the sections $\varphi_{j}$ by
$\varphi_{1}$ $=\chi_{1}\cdot e_{1,1},$ $\ldots,$ $\quad\varphi_{r}=\chi_{1}\cdot e_{1,r},$
$\quad\vdots$
$\varphi_{r(N-1)+1}$ $=\chi_{N}\cdot e_{N,1},$ $\ldots,$ $\quad\varphi_{rN}=\chi_{N}\cdot e_{N,r}.$
We compute, using Lemmas 2 and 3:
$\sum_{j=1}^{rN}|P\varphi_{j}|^{2}$ $=\sum_{j=1}^{N}\sum_{l=1}^{r}|P(\chi_{j}e_{j,l})|^{2}=\sum_{j=1}^{N}\chi_{j}^{2}\sum_{l=1}^{r}|Pe_{j,l}|^{2}=\sum_{j=1}^{N}\chi_{j}^{2}\cdot k=k.\qed$
4. Proof of the theorem
In this section we carry out the proof of the theorem stated in the introduction.
4.1. Definition of the constant C(M)
Let $M$ be an $n$-dimensional closed Riemannian spin manifold. We apply Proposition 2 to the spinor bundle $E=\Sigma M$ of $M$ if $n$ is even and to $E=\Sigma M\oplus\Sigma M$ if $n$ is odd. We rescale the resulting smooth sections $\varphi_{1},...,\varphi_{m}\in C^{\infty}(M,E)$ such that
$\sum_{j=1}^{m}|P\varphi_{j}|^{2}\equiv 1$
whenever $P\colon E\to F$ is the orthogonal projection onto a subbundle $F\subset E$ of half the rank of $E$.
Let $\tilde{D}_{M}$ be the intrinsic Dirac operator of $M$ and set $D_{M}:=\tilde{D}_{M}$ if $n$ is even and $D_{M}=\begin{pmatrix}\tilde{D}_{M}&0\\ 0&-\tilde{D}_{M}\end{pmatrix}$ if $n$ is odd. We define
$C(M):=4\sum_{j=1}^{m}\|D_{M}\varphi_{j}\|_{L^{2}(M)}\cdot\|\varphi_{j}\|_{L^{2}(M)}.$
4.2. The estimate
Now let $X$ be a spin fill-in of $M$, i.e., $M=\partial X$ as a Riemannian spin manifold. Without loss of generality, we may assume that $X$ is connected since otherwise we can apply the following argument to each connected component separately. Connected components of $X$ with empty boundary do not contribute to the integral of the mean curvature and can thus be ignored.
We identify $E=\Sigma X|_{M}$ as discussed in Section 2. Recall the splitting $E=\Sigma^{+}M\oplus\Sigma^{-}M$ in (2) and the orthogonal projections onto these subbundles $P^{\pm}\colon E\to\Sigma^{\pm}M$.
Let $\lambda\geq 0$ be such that $\mathrm{scal}_{X}\geq-\lambda^{2}$ on $X$. We apply Proposition 1 to $\varphi=P^{+}\varphi_{j}$ and $\Psi=0$ to obtain spinors $\Phi_{j}\in H^{1}(X,\Sigma X)$ with $(D-i\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2})\Phi_{j}=0$ and $P^{+}(\Phi_{j}|_{M})=P^{+}\varphi_{j}$. Since the boundary condition $P^{+}(\cdot|_{M})$ is $\infty$-regular, these spinors are smooth, $\Phi_{j}\in C^{\infty}(X,\Sigma X)$. The Weitzenböck formula and integration by parts give:
$0$ $=\int_{X}\Big\langle\Big(D+i\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\Big)\Big(D-i\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\Big)\Phi_{j},\Phi_{j}\Big\rangle$
$=\int_{X}\langle(D^{2}+\tfrac{n+1}{n}\tfrac{\lambda^{2}}{4})\Phi_{j},\Phi_{j}\rangle$
$=\int_{X}\langle(\nabla^{*}\nabla+\tfrac{1}{4}\mathrm{scal}_{X}+\tfrac{\lambda^{2}}{4}+\tfrac{1}{n}\tfrac{\lambda^{2}}{4})\Phi_{j},\Phi_{j}\rangle$
$\geq\int_{X}\langle(\nabla^{*}\nabla+\tfrac{1}{n}\tfrac{\lambda^{2}}{4})\Phi_{j},\Phi_{j}\rangle.$
We introduce a new connection $\tilde{\nabla}$ on $\Sigma X$ by setting
$\tilde{\nabla}_{X}\Phi:=\nabla_{X}\Phi+\tfrac{i}{\sqrt{n(n+1)}}\tfrac{\lambda}{2}\gamma(X)\Phi.$
One easily computes that
$\tilde{\nabla}^{*}\tilde{\nabla}=\nabla^{*}\nabla+\tfrac{1}{n}\tfrac{\lambda^{2}}{4}.$
It follows that
$0$ $\geq\int_{X}\langle\tilde{\nabla}^{*}\tilde{\nabla}\Phi_{j},\Phi_{j}\rangle$
$=\int_{X}|\tilde{\nabla}\Phi_{j}|^{2}+\int_{M}\langle\tilde{\nabla}_{\nu}\Phi_{j},\Phi_{j}\rangle$
$\geq\int_{M}\langle\tilde{\nabla}_{\nu}\Phi_{j},\Phi_{j}\rangle$
$=\int_{M}\langle\nabla_{\nu}\Phi_{j},\Phi_{j}\rangle+\frac{1}{\sqrt{n(n+1)}}\frac{\lambda}{2}\int_{M}\langle s\Phi_{j},\Phi_{j}\rangle.$
Using (3), we compute along the boundary:
$\nabla_{\nu}\Phi_{j}$ $=-\gamma(\nu)D\Phi_{j}-D_{M}\Phi_{j}+\tfrac{1}{2}H\Phi_{j}=i\,s\,D\Phi_{j}-D_{M}\Phi_{j}+\tfrac{1}{2}H\Phi_{j}=-\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\,s\,\Phi_{j}-D_{M}\Phi_{j}+\tfrac{1}{2}H\Phi_{j}.$
Hence, we find
$0$ $\geq\int_{M}\Big(-\sqrt{\tfrac{n+1}{n}}\tfrac{\lambda}{2}\langle s\Phi_{j},\Phi_{j}\rangle-\langle D_{M}\Phi_{j},\Phi_{j}\rangle+\tfrac{1}{2}H\,|\Phi_{j}|^{2}+\tfrac{1}{\sqrt{n(n+1)}}\tfrac{\lambda}{2}\langle s\Phi_{j},\Phi_{j}\rangle\Big)$
$=\int_{M}\Big(\sqrt{\tfrac{n}{n+1}}\tfrac{\lambda}{2}\langle P^{-}\Phi_{j}-P^{+}\Phi_{j},P^{+}\Phi_{j}+P^{-}\Phi_{j}\rangle-\langle D_{M}\Phi_{j},\Phi_{j}\rangle+\tfrac{1}{2}H\,|\Phi_{j}|^{2}\Big)$
$=\int_{M}\Big(\sqrt{\tfrac{n}{n+1}}\tfrac{\lambda}{2}(|P^{-}\Phi_{j}|^{2}-|P^{+}\Phi_{j}|^{2})-\langle D_{M}\Phi_{j},\Phi_{j}\rangle+\tfrac{1}{2}H\,|\Phi_{j}|^{2}\Big)$
$\geq\int_{M}\Big(-\sqrt{\tfrac{n}{n+1}}\tfrac{\lambda}{2}|P^{+}\Phi_{j}|^{2}-\langle D_{M}\Phi_{j},\Phi_{j}\rangle+\tfrac{1}{2}H\,|\Phi_{j}|^{2}\Big).$
Summation over $j=1,\ldots,m$ gives
$\frac{1}{2}\int_{M}H$ $=\frac{1}{2}\sum_{j=1}^{m}\int_{M}H|P^{+}\varphi_{j}|^{2}$
$=\frac{1}{2}\sum_{j=1}^{m}\int_{M}H|P^{+}\Phi_{j}|^{2}$
$=\frac{1}{2}\sum_{j=1}^{m}\int_{M}\Big[H|\Phi_{j}|^{2}-H|P^{-}\Phi_{j}|^{2}\Big]$
$\leq\sum_{j=1}^{m}\Big[\int_{M}\langle D_{M}\,\Phi_{j},\Phi_{j}\rangle+\sqrt{\frac{n}{n+1}}\frac{\lambda}{2}\int_{M}|P^{+}\Phi_{j}|^{2}-\frac{1}{2}\int_{M}H|P^{-}\Phi_{j}|^{2}\Big].$ (7)
For the first term we find
$\sum_{j=1}^{m}\int_{M}\langle D_{M}\,\Phi_{j},\Phi_{j}\rangle$ $=\sum_{j=1}^{m}\int_{M}(\langle D_{M}\,P^{+}\Phi_{j},P^{-}\Phi_{j}\rangle+\langle D_{M}\,P^{-}\Phi_{j},P^{+}\Phi_{j}\rangle)$
$=\sum_{j=1}^{m}\int_{M}(\langle D_{M}\,P^{+}\Phi_{j},P^{-}\Phi_{j}\rangle+\langle P^{-}\Phi_{j},D_{M}\,P^{+}\Phi_{j}\rangle)$
$=2\mathsf{Re}\sum_{j=1}^{m}\int_{M}\langle D_{M}\,P^{+}\Phi_{j},P^{-}\Phi_{j}\rangle$
$=2\mathsf{Re}\sum_{j=1}^{m}\int_{M}\langle D_{M}\,P^{+}\varphi_{j},P^{-}\Phi_{j}\rangle$
$=2\mathsf{Re}\sum_{j=1}^{m}\int_{M}\langle D_{M}\,\varphi_{j},P^{-}\Phi_{j}\rangle$
$\leq 2\sum_{j=1}^{m}\|D_{M}\varphi_{j}\|_{L^{2}(M)}\cdot\|P^{-}\Phi_{j}\|_{L^{2}(M)}$
$\leq 2\sum_{j=1}^{m}\|D_{M}\varphi_{j}\|_{L^{2}(M)}\cdot\|P^{+}\Phi_{j}\|_{L^{2}(M)}$ (8)
$=2\sum_{j=1}^{m}\|D_{M}\varphi_{j}\|_{L^{2}(M)}\cdot\|P^{+}\varphi_{j}\|_{L^{2}(M)}$
$\leq 2\sum_{j=1}^{m}\|D_{M}\varphi_{j}\|_{L^{2}(M)}\cdot\|\varphi_{j}\|_{L^{2}(M)}$
$=\tfrac{1}{2}C(M),$ (9)
where we applied Lemma 1 in (8). For the second term (7) we get
$\sum_{j=1}^{m}\sqrt{\frac{n}{n+1}}\frac{\lambda}{2}\int_{M}|P^{+}\Phi_{j}|^{2}$ $=\sqrt{\frac{n}{n+1}}\frac{\lambda}{2}\sum_{j=1}^{m}\int_{M}|P^{+}\varphi_{j}|^{2}=\sqrt{\frac{n}{n+1}}\frac{\lambda}{2}\mathsf{vol}(M).$ (10)
For the third term in (7) we obtain, again using Lemma 1:
$-\frac{1}{2}\sum_{j=1}^{m}\int_{M}H|P^{-}\Phi_{j}|^{2}$ $\leq\frac{\eta}{2}\sum_{j=1}^{m}\|P^{-}\Phi_{j}\|^{2}\leq\frac{\eta}{2}\sum_{j=1}^{m}\|P^{+}\Phi_{j}\|^{2}=\frac{\eta}{2}\sum_{j=1}^{m}\|P^{+}\varphi_{j}\|^{2}=\frac{\eta}{2}\mathsf{vol}(M).$ (11)
Combining (7), (9), (10), and (11) we find
$\int_{M}H\leq C(M)+\bigg(\sqrt{\frac{n}{n+1}}\lambda+\eta\bigg)\mathsf{vol}(M).$
This completes the proof of the theorem.
5. Concluding remarks
The dependence on the lower scalar curvature bound is very explicit and appears to be optimal as the following example shows:
Example.
For $\kappa\in\mathbb{R}$ let $\mathbb{M}_{\kappa}^{n+1}$ be the $(n+1)$-dimensional simply connected space form of constant sectional curvature $\kappa$. Let $X\subset\mathbb{M}_{\kappa}^{n+1}$ be a compact ball where the radius is chosen such that $M=\partial X$ is the standard unit sphere. This is possible if $\kappa\leq 1$. The mean curvature of $M$ is given by $H=n\sqrt{1-\kappa}$. For negative $\kappa$ we set $n(n+1)\kappa=-\lambda^{2}$ with $\lambda>0$. Then $\mathrm{scal}_{X}=-\lambda^{2}$ and we have
$\int_{M}H=n\cdot\sqrt{1+\tfrac{\lambda^{2}}{n(n+1)}}\cdot\mathsf{vol}(M)=\sqrt{\tfrac{n^{2}}{\lambda^{2}}+\tfrac{n}{n+1}}\cdot\lambda\cdot\mathsf{vol}(M).$
This shows that the dependence on $\lambda$ in our estimate is of the right order as $\lambda\to\infty$.
It seems strange that one needs a lower bound on the mean curvature in order to obtain an upper bound on its integral. It is unclear to the author whether this dependence can be removed. Note that the Gauss-Bonnet argument for $\dim(M)=1$ and $\lambda=0$ in the introduction does not need a lower bound on $H$.
Remark.
The definition of the constant $C(M)$ in Section 4.1 depends on the spin structure of $M$. However $M$, being compact, has only finitely many spin structures. Thus, $C(M)$ can be chosen independently of the spin structure.